A box holds 9 green marbles, 8 red marbles, and 7 blue marbles. Whats the probability of getting a green?
Answer by checkley75(3666)

9+8+7=24 total marbles in the box.
success/total=probability of a green marble
9/24=3/8=.37.5=37.5% probability of getting a green marble.